How many moles of PCl5 can be produced from 20.0 g of P4 (and excess Cl2)?
1
Expert's answer
2017-04-20T06:50:26-0400
Solution:

P4+10Cl2→4PCl5

n(P_4 )=(m(P_4))/(M(P_4))=20.0/124=0.16 mol
n(P_4 ):n(PCl_5 )=1:4

n(PCl_5 )=n(P_4 )×4=0.16×4=0.64 mol

Answer:
0.64 mol